OpenStudy (anonymous):

2^x+2^y=2^z - x, y, z values must be integers and not powers of 10; prove whether this is true or false for x,y,z values >2.

OpenStudy (anonymous):

So the thing I'm trying to prove is: \[\exists( x,y,z) \subset \mathbb{Z}^+: (x,y,z >2),( x,y,z \neq10), (2^x+2^y = 2^z)\]

OpenStudy (anonymous):

An example of a specific x,y, and z will do. Try a = 3, b = 3, c = 4

OpenStudy (anonymous):

confirm x,y,z >2? confirm x,y,z != 10? confirm 2^x + 2^y = 2^z?

OpenStudy (anonymous):

My answer's no good, eh?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

good try my man

OpenStudy (anonymous):

What's wrong with it? Are x and y not allowed to be equal?

OpenStudy (anonymous):

they can be

OpenStudy (anonymous):

Well then you get infinitely many solutions. Any x,y,z such that x=y z = x+1

OpenStudy (anonymous):

Oh, and x,z != 10.

OpenStudy (anonymous):

oh damn, hang on

OpenStudy (anonymous):

sorry wrong question lol

OpenStudy (anonymous):

i'll retype it

OpenStudy (anonymous):

3,3,4 is just one example. 4,4,5 is another.